### v7

@@ 1 Introduction
Transfer learning (TL) is a common practice for medical image classification (MIC), especially when training data are limited. In typical TL pipelines for MIC, deep convolutional neural networks (DCNNs) pretrained on large-scale source tasks (e.g., object recognition on ImageNet (Deng et al.(2009)Deng, Dong, Socher, Li, Kai Li, and Li Fei-Fei)) are finetuned as backbone models for target MIC tasks; see, e.g., Shin et al.(2016)Shin, Roth, Gao, Lu, Xu, Nogues, Yao, Mollura, and Summers; van Tulder and de Bruijne(2016); Rehman et al.(2020)Rehman, Naz, Razzak, Akram, and Imran; Huynh et al.(2016)Huynh, Li, and Giger; Antropova et al.(2017)Antropova, Huynh, and Giger; Ghafoorian et al.(2017)Ghafoorian, Mehrtash, Kapur, Karssemeijer, Marchiori, Pesteie, Guttmann, de Leeuw, Tempany, Van Ginneken, et al., for examples of prior successes.
@@
The key to TL is feature reuse from the source to the target tasks, which leads to practical benefits such as fast convergence in training, and good test performance even if the target data are scarce (Yosinski et al.(2014)Yosinski, Clune, Bengio, and Lipson). Pretrained DCNNs extract increasingly more abstract visual features from bottom to top layers: from low-level corners and textures, to mid-level blobs and parts, and finally to high-level shapes and patterns (Zeiler and Fergus(2014); Yosinski et al.(2014)Yosinski, Clune, Bengio, and Lipson). While shapes and patterns are crucial for recognizing and segmenting generic visual objects (see Fig. 1 (left)), they are not necessarily the defining features for diseases: diseases can often take the form of abnormal textures and blobs, which correspond to low- to mid-level features (see Fig. 1 (right)). So intuitively for MIC, we may only need to finetune a reasonable number of the bottom layers commensurate with the levels of features needed, and ignore the top layers. However, standard TL practice for MIC retains all layers, and uses them as fixed feature extractors or finetunes them uniformly.
@@
Tajbakhsh et al.(2016)Tajbakhsh, Shin, Gurudu, Hurst, Kendall, Gotway, and Liang; Raghu et al.(2019b)Raghu, Zhang, Kleinberg, and Bengio depart from the uniform TL approach and propose TL methods that treat top and bottom layers differently. Prioritizing high-level features and the classifier, Tajbakhsh et al.(2016)Tajbakhsh, Shin, Gurudu, Hurst, Kendall, Gotway, and Liang proposes layer-wise finetuning (LWFT) that finetunes an appropriate number of top layers and freezes the remaining bottom layers. In comparison, to improve training speed while preserving performance, Raghu et al.(2019b)Raghu, Zhang, Kleinberg, and Bengio proposes TransFusion (TF) that finetunes bottom layers but retrains a coarsened version of top layers from scratch.
@@
Figure 1: (left) The feature hierarchy learned by typical DCNNs, see Appendix E for details; (right) Examples of diseases in a chest x-ray Nguyen et al.(2021)Nguyen, Pham, Nguyen, Nguyen, Dao, Vu, Lam, and Le.
@@
Neither of the above differential TL strategies clearly address the conundrum of why top layers are needed when only the features in bottom layers are to be reused. To bridge the gap, in this paper, we propose a novel, perhaps radical, TL strategy: remove top layers after an appropriate cutoff point, and finetune the truncated model left, dubbed TruncatedTL (TTL)—this is entirely consistent with our intuition about the feature hierarchy. Our main contributions include: 1) confirming the deficiency of full TL. By experimenting with full and differential TL strategies—including our TTL—on three MIC tasks and three popularly used DCNN models for MIC, we find that full TL (FTL) is almost always suboptimal in terms of classification performance, confirming the observation in Tajbakhsh et al.(2016)Tajbakhsh, Shin, Gurudu, Hurst, Kendall, Gotway, and Liang. 2) proposing TruncatedTL (TTL) that leads to effective and compact models. Our TTL outperforms other differential TL methods, while the resulting models are always smaller, sometimes substantially so. This leads to reduced computation and fast prediction during inference, and can be particularly valuable when dealing with 3D medical data such as CT and MRI images. 3) quantifying feature transferability in TL for MIC. We use singular vector canonical correlation analysis (SVCCA) (Raghu et al.(2017)Raghu, Gilmer, Yosinski, and Sohl-Dickstein) to analyze feature transferability and confirm the importance of low- and mid-level features for MIC. The quantitative analysis also provides insights on how to choose high-quality truncation points to further optimize the model performance and efficiency of TTL.
@@ 2 Related Work
Deep TL TL by reusing and finetuning visual features learned in DCNNs entered computer vision (CV) once DCNNs became the cornerstone of state-of-the-art (SOTA) object recognition models back to 2012. For example, Donahue et al.(2014)Donahue, Jia, Vinyals, Hoffman, Zhang, Tzeng, and Darrell; Sharif Razavian et al.(2014)Sharif Razavian, Azizpour, Sullivan, and Carlsson propose using part of or full pretrained DCNNs as feature extractors for generic visual recognition. Yosinski et al.(2014)Yosinski, Clune, Bengio, and Lipson studies the hierarchy of DCNN-based visual features, characterizes their transferability, and proposes finetuning pretrained features to boost performance on target tasks. Moreover, Erhan et al.(2009)Erhan, Bengio, Courville, and Vincent; Zeiler and Fergus(2014) propose techniques to visualize visual features and their hierarchy. This popular line of TL techniques is among the broad family of TL methods for knowledge transfer from source to target tasks based on deep networks (Tan et al.(2018)Tan, Sun, Kong, Zhang, Yang, and Liu) and other learning models (Zhuang et al.(2021)Zhuang, Qi, Duan, Xi, Zhu, Zhu, Xiong, and He), and is the focus of this paper.
@@
General TL stategies While pretrained DCNNs can be either used as fixed feature extractors, or partially or fully finetuned on the target data, Yosinski et al.(2014)Yosinski, Clune, Bengio, and Lipson argues that bottom layers learn generic features while top layers learn task-specific features, leading to folklore guidelines on how to choose appropriate TL strategies in various scenarios, as summarized in Fig. 2.
@@
Our intuition about the feature hierarchy is slightly different: bottom layers learn low-level features that are spatially localized, and top layers learn high-level features that are spatially extensive (see Fig. 1). Although the two hierarchies may be aligned for most cases, they are distinguished by whether spatial scales are considered: task-specific features may be spatially localized, e.g., texture features to classify skin lesions (Tschandl et al.(2018)Tschandl, Rosendahl, and Kittler), and general features may be spatially extensive, e.g., generic brain silhouettes in brain MRI images. Moreover, the spatial-scale hierarchy is built into DCNNs by design (Bronstein et al.(2021)Bronstein, Bruna, Cohen, and Veličković). So we argue that the intuition about the low-high spatial feature hierarchy is more pertinent. We note that none of the popular strategies as summarized in Fig. 2 modify the pretrained DCNN models (except for the final multi-layer perceptron, MLP, classifiers)—in contrast to TF and our TTL.
@@
Differential TL Early work on TL for MIC (Shin et al.(2016)Shin, Roth, Gao, Lu, Xu, Nogues, Yao, Mollura, and Summers; van Tulder and de Bruijne(2016); Rehman et al.(2020)Rehman, Naz, Razzak, Akram, and Imran; Huynh et al.(2016)Huynh, Li, and Giger; Antropova et al.(2017)Antropova, Huynh, and Giger; Hosny et al.(2018)Hosny, Kassem, and Foaud; Sun et al.(2021)Sun, Peng, Li, Adila, Zaiman, Melton, Ingraham, Murray, Boley, Switzer, et al.) and segmentation (Van Opbroek et al.(2014)Van Opbroek, Ikram, Vernooij, and De Bruijne; Ronneberger et al.(2015)Ronneberger, Fischer, and Brox; Ghafoorian et al.(2017)Ghafoorian, Mehrtash, Kapur, Karssemeijer, Marchiori, Pesteie, Guttmann, de Leeuw, Tempany, Van Ginneken, et al.; Jiang et al.(2018)Jiang, Zhang, Wang, and Ko; Karimi et al.(2021)Karimi, Warfield, and Gholipour) parallels the relevant developments in CV, and mostly uses DCNNs pretrained on CV tasks as feature extractors or initializers (i.e., for finetuning). In fact, these two strategies remain dominant according to the very recent survey (Kim et al.(2022)Kim, Cosa-Linan, Santhanam, Jannesari, Maros, and Ganslandt) on TL for MIC, which reviews around $120$ relevant papers. But the former seems inappropriate as medical data are disparate from natural images dealt with in CV. From the bottom panel of Fig. 2, finetuning at least part of the DCNNs is probably more competitive even if the target data are limited. In this line, LWFT (Tajbakhsh et al.(2016)Tajbakhsh, Shin, Gurudu, Hurst, Kendall, Gotway, and Liang) finetunes top layers and freeze bottom layers, and incrementally allows finetuning more layers during model selection—this is recommended in Kim et al.(2022)Kim, Cosa-Linan, Santhanam, Jannesari, Maros, and Ganslandt as a practical TL strategy for MIC that strikes a balance between training efficiency and performance. Similarly, TF (Raghu et al.(2019b)Raghu, Zhang, Kleinberg, and Bengio) coarsens top layers which are then trained from scratch, and finetunes bottom layers from pretrained weights. Both LWFT and TF take inspiration from the general-specific feature hierarchy. In contrast, motivated by the low-high spatial feature hierarchy, our novel TTL method removes the top layers entirely and directly finetunes the truncated models. Our experiments in Section 4 confirm that TTL surpasses LWFT and TF with improved performance and reduced inference cost.
@@
Compact models for MIC Both TF and our TTL lead to reduced models that can boost the inference efficiency, the first of its kind in TL for MIC, although the TF paper (Raghu et al.(2019b)Raghu, Zhang, Kleinberg, and Bengio) does not stress this point. Compact models have been designed for specific MIC tasks, e.g., Raghu et al.(2019b)Raghu, Zhang, Kleinberg, and Bengio; Huang et al.(2020b)Huang, Kothari, Banerjee, Chute, Ball, Borus, Huang, Patel, Rajpurkar, Irvin, et al., but our evaluation on the task of Huang et al.(2020b)Huang, Kothari, Banerjee, Chute, Ball, Borus, Huang, Patel, Rajpurkar, Irvin, et al. in Section 4 suggests that the differential TL strategies based on generic pretrained models, particularly our TTL, can outperform TL based on handcrafted models. Moreover, the growing set of methods for model quantization and compression (Wang et al.(2020)Wang, Chen, He, and Cheng; Krishnamoorthi(2018); Polino et al.(2018)Polino, Pascanu, and Alistarh; Han et al.(2015)Han, Mao, and Dally) are equally applicable to both the original models and the reduced models.
@@ 3 Efficient transfer learning for MIC
Let $\mathcal{X}\times\mathcal{Y}$ denote any input-output (or feature-label) product space, and $\mathcal{D}_{\mathcal{X}\times\mathcal{Y}}$ a distribution on $\mathcal{X}\times\mathcal{Y}$. TL considers a source task $\mathcal{D}_{\mathcal{X}_{s}\times\mathcal{Y}_{s}}\mapsto p_{s}$, where $p_{s}$ is a desired predictor, and a target task $\mathcal{D}_{\mathcal{X}_{t}\times\mathcal{Y}_{t}}\mapsto p_{t}$. In typical TL, $\mathcal{X}_{t}\times\mathcal{Y}_{t}$ may be different from $\mathcal{X}_{s}\times\mathcal{Y}_{s}$, or at least $\mathcal{D}_{\mathcal{X}_{t}\times\mathcal{Y}_{t}}\neq\mathcal{D}_{\mathcal{X}_{s}\times\mathcal{Y}_{s}}$ even if $\mathcal{X}_{t}\times\mathcal{Y}_{t}=\mathcal{X}_{s}\times\mathcal{Y}_{s}$. The goal of TL is to transfer the knowledge from the source task $\mathcal{D}_{\mathcal{X}_{s}\times\mathcal{Y}_{s}}\mapsto p_{s}$ that is solved beforehand to the target task $\mathcal{D}_{\mathcal{X}_{t}\times\mathcal{Y}_{t}}\mapsto p_{t}$ (Pan and Yang(2009); Tan et al.(2018)Tan, Sun, Kong, Zhang, Yang, and Liu).
@@ 3.1 Prior differential TL approaches
The two differential TL methods for MIC, i.e., LWFT (Tajbakhsh et al.(2016)Tajbakhsh, Shin, Gurudu, Hurst, Kendall, Gotway, and Liang) and TF (Raghu et al.(2019a)Raghu, Blumer, Sayres, Obermeyer, Kleinberg, Mullainathan, and Kleinberg), differ from the dominant TL approaches in that they treat the top and bottom layers differently, as illustrated in Fig. 3 (i).
Layer-wise finetuning LWFT does not distinguish MLP layers and convolutional layers. So slightly abusing our notation, assume that the pretrained DCNN is $f_{N}\circ f_{N-1}\circ\dots\circ f_{1}$, where $N$ is the total number of layers including both the MLP and convolutional layers. LWFT finetunes the top $k$ layers $f_{N}\circ f_{N-1}\circ\dots\circ f_{N-k+1}$, and freezes the bottom $N-k$ layers $f_{N-k}\circ\dots\circ f_{1}$. The top layer $f_{N}$ is finetuned with a base learning rate (LR) $\eta$, and the other $k-1$ layers with a LR $\eta/10$. To find an appropriate $k$, Tajbakhsh et al.(2016)Tajbakhsh, Shin, Gurudu, Hurst, Kendall, Gotway, and Liang proposes an incremental model selection procedure: start with $k=1$ layer, and include one more layer into finetuning if the previous set of layers does not achieve the desired level of performance [^1] Finetuning starts with the original pretrained weights each time. . Although LWFT was originally proposed for AlexNet (Krizhevsky et al.(2012)Krizhevsky, Sutskever, and Hinton), it can be generalized to work with advanced DCNN models such as ResNets and DenseNets that have block structures.
@@ 3.2 Our truncated TL approach
To quickly confirm the efficiency of TTL, we apply TTL and competing TL methods on an X-ray-based COVID-19 classification task on the BIMCV dataset (Vayá et al.(2020)Vayá, Saborit, Montell, Pertusa, Bustos, Cazorla, Galant, Barber, Orozco-Beltrán, García-García, et al.); more details about the setup can be found in Section 4. We pick the COVID example, as the salient radiological patterns in COVID X-rays such as multifocal and bilateral ground glass opacities and consolidations are low- to mid-level visual features (Shi et al.(2020)Shi, Han, Jiang, Cao, Alwalid, Gu, Fan, and Zheng) and hence we can easily see the benefit of differential TL methods including our TTL. We measure the classification performance by both AUROC (area under the receiver-operating-characteristic curve) and AUPRC (area under the precision-recall curve), and measure the inference complexity by P (number of parameters in the model, M—millions), MACs (multiply-add operation counts (Sovrasov(2019)) [^2] https://github.com/sovrasov/flops-counter.pytorch, G—billion), T (wallclock run time in milliseconds; details of our computing environment can be found supplementary material).

### v8

@@ 1 Introduction
Transfer learning (TL) is a common practice for medical image classification (MIC), especially when training data are limited. In typical TL pipelines for MIC, deep convolutional neural networks (DCNNs) pretrained on large-scale source tasks (e.g., object recognition on ImageNet (Deng et al., 2009)) are finetuned as backbone models for target MIC tasks; see, e.g., Shin et al. (2016); van Tulder and de Bruijne (2016); Rehman et al. (2020); Huynh et al. (2016); Antropova et al. (2017); Ghafoorian et al. (2017), for examples of prior successes.
@@
The key to TL is feature reuse from the source to the target tasks, which leads to practical benefits such as fast convergence in training, and good test performance even if the target data are scarce (Yosinski et al., 2014). Pretrained DCNNs extract increasingly more abstract visual features from bottom to top layers: from low-level corners and textures, to mid-level blobs and parts, and finally to high-level shapes and patterns (Zeiler and Fergus, 2014; Yosinski et al., 2014). While shapes and patterns are crucial for recognizing and segmenting generic visual objects (see Fig. 1 (left)), they are not necessarily the defining features for diseases: diseases can often take the form of abnormal textures and blobs, which correspond to low- to mid-level features (see Fig. 1 (right)). So intuitively for MIC, we may only need to finetune a reasonable number of the bottom layers commensurate with the levels of features needed, and ignore the top layers. However, standard TL practice for MIC retains all layers, and uses them as fixed feature extractors or finetunes them uniformly.
@@
Tajbakhsh et al. (2016); Raghu et al. (2019b) depart from the uniform TL approach and propose TL methods that treat top and bottom layers differently. Prioritizing high-level features and the classifier, Tajbakhsh et al. (2016) proposes layer-wise finetuning (LWFT) that finetunes an appropriate number of top layers and freezes the remaining bottom layers. In comparison, to improve training speed while preserving performance, Raghu et al. (2019b) proposes TransFusion (TF) that finetunes bottom layers but retrains a coarsened version of top layers from scratch.
@@
Figure 1: (left) The feature hierarchy learned by typical DCNNs, see Appendix E for details; (right) Examples of diseases in a chest x-ray Nguyen et al. (2021).
@@
Neither of the above differential TL strategies clearly address the conundrum of why top layers are needed when only the features in bottom layers are to be reused. To bridge the gap, in this paper, we propose a novel, perhaps radical, TL strategy: remove top layers after an appropriate cutoff point, and finetune the truncated model left, dubbed TruncatedTL (TTL)—this is entirely consistent with our intuition about the feature hierarchy. Our main contributions include: 1) confirming the deficiency of full TL. By experimenting with full and differential TL strategies—including our TTL—on three MIC tasks and three popularly used DCNN models for MIC, we find that full TL (FTL) is almost always suboptimal in terms of classification performance, confirming the observation in Tajbakhsh et al. (2016). 2) proposing TruncatedTL (TTL) that leads to effective and compact models. Our TTL outperforms other differential TL methods, while the resulting models are always smaller, sometimes substantially so. This leads to reduced computation and fast prediction during inference, and can be particularly valuable when dealing with 3D medical data such as CT and MRI images. 3) quantifying feature transferability in TL for MIC. We use singular vector canonical correlation analysis (SVCCA) (Raghu et al., 2017) to analyze feature transferability and confirm the importance of low- and mid-level features for MIC. The quantitative analysis also provides insights on how to choose high-quality truncation points to further optimize the model performance and efficiency of TTL.
@@ 2 Related Work
Deep TL TL by reusing and finetuning visual features learned in DCNNs entered computer vision (CV) once DCNNs became the cornerstone of state-of-the-art (SOTA) object recognition models back to 2012. For example, Donahue et al. (2014); Sharif Razavian et al. (2014) propose using part of or full pretrained DCNNs as feature extractors for generic visual recognition. Yosinski et al. (2014) studies the hierarchy of DCNN-based visual features, characterizes their transferability, and proposes finetuning pretrained features to boost performance on target tasks. Moreover, Erhan et al. (2009); Zeiler and Fergus (2014) propose techniques to visualize visual features and their hierarchy. This popular line of TL techniques is among the broad family of TL methods for knowledge transfer from source to target tasks based on deep networks (Tan et al., 2018) and other learning models (Zhuang et al., 2021), and is the focus of this paper.
@@
General TL stategies While pretrained DCNNs can be either used as fixed feature extractors, or partially or fully finetuned on the target data, Yosinski et al. (2014) argues that bottom layers learn generic features while top layers learn task-specific features, leading to folklore guidelines on how to choose appropriate TL strategies in various scenarios, as summarized in Fig. 2.
@@
Our intuition about the feature hierarchy is slightly different: bottom layers learn low-level features that are spatially localized, and top layers learn high-level features that are spatially extensive (see Fig. 1). Although the two hierarchies may be aligned for most cases, they are distinguished by whether spatial scales are considered: task-specific features may be spatially localized, e.g., texture features to classify skin lesions (Tschandl et al., 2018), and general features may be spatially extensive, e.g., generic brain silhouettes in brain MRI images. Moreover, the spatial-scale hierarchy is built into DCNNs by design (Bronstein et al., 2021). So we argue that the intuition about the low-high spatial feature hierarchy is more pertinent. We note that none of the popular strategies as summarized in Fig. 2 modify the pretrained DCNN models (except for the final multi-layer perceptron, MLP, classifiers)—in contrast to TF and our TTL.
@@
Differential TL Early work on TL for MIC (Shin et al., 2016; van Tulder and de Bruijne, 2016; Rehman et al., 2020; Huynh et al., 2016; Antropova et al., 2017; Hosny et al., 2018; Sun et al., 2021) and segmentation (Van Opbroek et al., 2014; Ronneberger et al., 2015; Ghafoorian et al., 2017; Jiang et al., 2018; Karimi et al., 2021) parallels the relevant developments in CV, and mostly uses DCNNs pretrained on CV tasks as feature extractors or initializers (i.e., for finetuning). In fact, these two strategies remain dominant according to the very recent survey (Kim et al., 2022) on TL for MIC, which reviews around $120$ relevant papers. But the former seems inappropriate as medical data are disparate from natural images dealt with in CV. From the bottom panel of Fig. 2, finetuning at least part of the DCNNs is probably more competitive even if the target data are limited. In this line, LWFT (Tajbakhsh et al., 2016) finetunes top layers and freeze bottom layers, and incrementally allows finetuning more layers during model selection—this is recommended in Kim et al. (2022) as a practical TL strategy for MIC that strikes a balance between training efficiency and performance. Similarly, TF (Raghu et al., 2019b) coarsens top layers which are then trained from scratch, and finetunes bottom layers from pretrained weights. Both LWFT and TF take inspiration from the general-specific feature hierarchy. In contrast, motivated by the low-high spatial feature hierarchy, our novel TTL method removes the top layers entirely and directly finetunes the truncated models. Our experiments in Section 4 confirm that TTL surpasses LWFT and TF with improved performance and reduced inference cost.
@@
Compact models for MIC Both TF and our TTL lead to reduced models that can boost the inference efficiency, the first of its kind in TL for MIC, although the TF paper (Raghu et al., 2019b) does not stress this point. Compact models have been designed for specific MIC tasks, e.g., Raghu et al. (2019b); Huang et al. (2020b), but our evaluation on the task of Huang et al. (2020b) in Section 4 suggests that the differential TL strategies based on generic pretrained models, particularly our TTL, can outperform TL based on handcrafted models. Moreover, the growing set of methods for model quantization and compression (Wang et al., 2020; Krishnamoorthi, 2018; Polino et al., 2018; Han et al., 2015) are equally applicable to both the original models and the reduced models.
@@ 3 Efficient transfer learning for MIC
Let $\mathcal{X}\times\mathcal{Y}$ denote any input-output (or feature-label) product space, and $\mathcal{D}_{\mathcal{X}\times\mathcal{Y}}$ a distribution on $\mathcal{X}\times\mathcal{Y}$. TL considers a source task $\mathcal{D}_{\mathcal{X}_{s}\times\mathcal{Y}_{s}}\mapsto p_{s}$, where $p_{s}$ is a desired predictor, and a target task $\mathcal{D}_{\mathcal{X}_{t}\times\mathcal{Y}_{t}}\mapsto p_{t}$. In typical TL, $\mathcal{X}_{t}\times\mathcal{Y}_{t}$ may be different from $\mathcal{X}_{s}\times\mathcal{Y}_{s}$, or at least $\mathcal{D}_{\mathcal{X}_{t}\times\mathcal{Y}_{t}}\neq\mathcal{D}_{\mathcal{X}_{s}\times\mathcal{Y}_{s}}$ even if $\mathcal{X}_{t}\times\mathcal{Y}_{t}=\mathcal{X}_{s}\times\mathcal{Y}_{s}$. The goal of TL is to transfer the knowledge from the source task $\mathcal{D}_{\mathcal{X}_{s}\times\mathcal{Y}_{s}}\mapsto p_{s}$ that is solved beforehand to the target task $\mathcal{D}_{\mathcal{X}_{t}\times\mathcal{Y}_{t}}\mapsto p_{t}$ (Pan and Yang, 2009; Tan et al., 2018).
@@ 3.1 Prior differential TL approaches
The two differential TL methods for MIC, i.e., LWFT (Tajbakhsh et al., 2016) and TF (Raghu et al., 2019a), differ from the dominant TL approaches in that they treat the top and bottom layers differently, as illustrated in Fig. 3 (i).
Layer-wise finetuning LWFT does not distinguish MLP layers and convolutional layers. So slightly abusing our notation, assume that the pretrained DCNN is $f_{N}\circ f_{N-1}\circ\dots\circ f_{1}$, where $N$ is the total number of layers including both the MLP and convolutional layers. LWFT finetunes the top $k$ layers $f_{N}\circ f_{N-1}\circ\dots\circ f_{N-k+1}$, and freezes the bottom $N-k$ layers $f_{N-k}\circ\dots\circ f_{1}$. The top layer $f_{N}$ is finetuned with a base learning rate (LR) $\eta$, and the other $k-1$ layers with a LR $\eta/10$. To find an appropriate $k$, Tajbakhsh et al. (2016) proposes an incremental model selection procedure: start with $k=1$ layer, and include one more layer into finetuning if the previous set of layers does not achieve the desired level of performance [^1] Finetuning starts with the original pretrained weights each time. . Although LWFT was originally proposed for AlexNet (Krizhevsky et al., 2012), it can be generalized to work with advanced DCNN models such as ResNets and DenseNets that have block structures.
@@ 3.2 Our truncated TL approach
To quickly confirm the efficiency of TTL, we apply TTL and competing TL methods on an X-ray-based COVID-19 classification task on the BIMCV dataset (Vayá et al., 2020); more details about the setup can be found in Section 4. We pick the COVID example, as the salient radiological patterns in COVID X-rays such as multifocal and bilateral ground glass opacities and consolidations are low- to mid-level visual features (Shi et al., 2020) and hence we can easily see the benefit of differential TL methods including our TTL. We measure the classification performance by both AUROC (area under the receiver-operating-characteristic curve) and AUPRC (area under the precision-recall curve), and measure the inference complexity by P (number of parameters in the model, M—millions), MACs (multiply-add operation counts (Sovrasov, 2019) [^2] https://github.com/sovrasov/flops-counter.pytorch, G—billion), T (wallclock run time in milliseconds; details of our computing environment can be found supplementary material).